OpenStudy (anonymous):

prove d/dx (1/u^n)= -n/u^n+1 * du/dx

OpenStudy (anonymous):

\[\frac{1}{u^n}=u^{-n}\] does that help?

OpenStudy (anonymous):

chain rule here i guess. since we are assuming that \(u=u(x)\) then \(y=u^{-n}\) and \(\frac{dy}{du}=-nu^{-n-1}\) \[\frac{dy}{dx}=\frac{dy}{du}\times \frac{du}{dx}\]

OpenStudy (anonymous):

@satellite73 how about this one prove d/dx (square root of u) = 1/2*square root of u * du/dx?? help me pls..

OpenStudy (anonymous):

thank you..

OpenStudy (helder_edwin):

it is pretty much as Satellite did before \(u=u(x)\) and \(y=\sqrt{u}\) then \[ \large \frac{dy}{dx}=\frac{dy}{du}\cdot\frac{du}{dx} \] where \[ \large \frac{dy}{du}=\frac{1}{2\sqrt{u}} \]

OpenStudy (anonymous):

where did u get u=u(x)?

OpenStudy (helder_edwin):

it is a way of saying that \(u\) is a dependent variable and that it depends on \(x\)

OpenStudy (helder_edwin):

your post are about the chain rule as Satellite pointed out.

OpenStudy (anonymous):

yeah..right.. thank you..

OpenStudy (helder_edwin):

did u get it?

OpenStudy (anonymous):

yes i got it. thanks.

OpenStudy (helder_edwin):

u r welcome.